OpenStudy (anonymous):

An apple pie has a diameter of 8 inches. What is the area of the pie?

OpenStudy (btaylor):

Use these formulas: \(diameter = 2 \times radius\) and \(Area=\pi r^2\)

OpenStudy (anonymous):

First find the radius and to find the radius u have to divide diameter by 2 which is 8/2=4/1=4 Now use the formula of area of a circle which is pie x r square Which will be Pie (4) square Which is 50.265 After rounding off The answer will be 50.3 or 50

OpenStudy (anonymous):

Sorry 50.3 inches square or 50 inches square